OpenStudy (jenniferjuice):

Calculate the average rate of change for the function f(x) = −x4 + 4x3 − 2x2 + x + 1, from x = 0 to x = 1. 0 1 2 7

myininaya (myininaya):

to find the average rate use the formula to find the slope of the secant line going though f from x=0 to x=1. that is \[\frac{f(1)-f(0)}{1-0}\]

OpenStudy (jenniferjuice):

f1/1?

myininaya (myininaya):

f(0)=?

OpenStudy (jenniferjuice):

is my answer 1 ?

myininaya (myininaya):

what is f(1)? what is f(0)?

OpenStudy (triciaal):

find f(x) when x = 1 find f(x) when x = 0

OpenStudy (triciaal):

isn't the rate of change the derivative of the function?

myininaya (myininaya):

yep but it says average rate of change

myininaya (myininaya):

which just calls for the slope formula without all the limit stuff

OpenStudy (jenniferjuice):

so what do i do..?

myininaya (myininaya):

triciaal said you need to find f(x) when x=1 and also for x=0

myininaya (myininaya):

then plug into the above formula i mentioned

OpenStudy (triciaal):

@myininaya is your answer 7?

myininaya (myininaya):

Let's go in steps @jenniferjuice Can you find f(1)?

myininaya (myininaya):

you said f(x)=-x^4+4x^3-2x^2+x+1 you should be able to replace all the x's there with 1's to find f(1)

myininaya (myininaya):

like so f(1)=-1^4+4(1)^3-2(1)^2+1+1

myininaya (myininaya):

now simplify that.

OpenStudy (jenniferjuice):

how?

myininaya (myininaya):

what is 1^4?

myininaya (myininaya):

\[f(1)=-1^4+4(1)^3-2(1)^2+1+1 \\ -1+4(1)^3-2(1)^2+1+1\] so above i replaced 1^4 with 1 you should know what 1^3 and 1^2 is...

myininaya (myininaya):

1^4=1*1*1*1=1 1^3=1*1*1=? 1^2=1*1=? 1 to any power equals?

OpenStudy (jenniferjuice):

1

myininaya (myininaya):

right so replace the 1^n 's with 1

OpenStudy (jenniferjuice):

-1+4-2+2

myininaya (myininaya):

\[f(1)=-1^4+4(1)^3-2(1)^2+1+1 \\ -1 +4(1)-2(1)+1+1\] right -1+4-2+2 this can be simplified more

OpenStudy (jenniferjuice):

-1+4-2+2 -1+4-4 3-4 -1

myininaya (myininaya):

-2+2 isn't 4 or -4

myininaya (myininaya):

if you a credit of 2 then a debit of 2 then you have?

OpenStudy (jenniferjuice):

0

myininaya (myininaya):

so -2+2 is 0

myininaya (myininaya):

so -1+4 =?

OpenStudy (jenniferjuice):

3

myininaya (myininaya):

right so f(1) is 3 what is f(0)?

myininaya (myininaya):

f(0) should be the easiest to find

OpenStudy (jenniferjuice):

is it 0

myininaya (myininaya):

nope

myininaya (myininaya):

remember f(0) tells you to use the machine (function ) that is labeled f

myininaya (myininaya):

f(x)=-x^4+4x^3-2x^2+x+1

myininaya (myininaya):

to find f(0) replace every single x there with 0 then simplify

myininaya (myininaya):

look if you replace all of those x's with 0 won't you just have 0+0-0+0+1?

OpenStudy (jenniferjuice):

yes so 1

OpenStudy (jenniferjuice):

@triciaal help here?

OpenStudy (triciaal):

for f(1) you got 3 for f(0) you got 1 using the formula she originally gave f(1) - (f 0)divided by 1-0 it would be (3-1)/1 = 2

OpenStudy (triciaal):

thanks for the medal